OpenStudy (anonymous):

Use properties of logarithms to condense the logarithmic expression. Write the expression as a single logarithm whose coefficient is 1. Evaluate logarithmic expressions if possible. 4lnx-1/5lny = ?

jhonyy9 (jhonyy9):

do you know the proprties of logarithms ?

jhonyy9 (jhonyy9):

loga +logb= ?

jhonyy9 (jhonyy9):

and loga -logb= ?

jhonyy9 (jhonyy9):

and 2loga +3logb= ?

jhonyy9 (jhonyy9):

so loga +logb=log(a*b) loga-logb=log(a/b) 2loga +3logb=loga^2 +logb^3 =log((a^2)*(b^3)) ok ?

OpenStudy (anonymous):

Yes, that makes sense

OpenStudy (anonymous):

So would it be... ln((x^4/lny^1/5)) ??

jhonyy9 (jhonyy9):

how you have got it ?

jhonyy9 (jhonyy9):

so if there are 4lnx so this will be lnx^4 yes ?

jhonyy9 (jhonyy9):

and 1/5lny =lny^(1/5) yes ?

jhonyy9 (jhonyy9):

and you know again that lna -lnb=ln(a/b)

jhonyy9 (jhonyy9):

so than what will be the right answer ?

jhonyy9 (jhonyy9):

came on ,please !!!

jhonyy9 (jhonyy9):

your answer is aproximatly right just can you tel me why you have wrote the ln on denominator too ?

OpenStudy (anonymous):

I was just trying to apply the rules you mentioned. Would it just be ln((x^4/y^1/5))?

OpenStudy (anonymous):

Sorry, trying to write down all the differant rules for notes...

jhonyy9 (jhonyy9):

yes ,right sure because lna -lnb not is equal lna/lnb ... so this is equal ln(a/b) ok ?

OpenStudy (anonymous):

Yes...ty

jhonyy9 (jhonyy9):

yw was my pleasure good luck bye